OpenStudy (anonymous):

In how many ways can we distribute 10 identical looking pencils to 4 students so that each student gets at least 1 pencil?

OpenStudy (ilovecake):

210 is your answer yw

OpenStudy (anonymous):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (paxpolaris):

the 1st student can get get upto 7 pencils: if she/he gets 7: 1 for everyone else \(\large1\) way to do that. if she/he gets 6: 2pencils for 1 student, 1 each for the other 2. \(\large3\) ways to do that. if 1st student gets 5: 3 for 1 student and 1 for other 2 (3+1+1). OR 2 for two students & 1 for one (2+2+1) \(3+3=\large6\) ways to do that. if 1st student gets 4: the others get (4+1+1) OR (3+2+1) OR (2+2+2) \(3+6+1=\large{10}\) ways to do that. if 1st student gets 3: the others get (5+1+1) OR (4+2+1) OR (3+2+2) OR (3+3+1) \(3+6+3+3=\large{15}\) ways to do that. if 1st student gets 2: the others get (7+1+1) OR (6+2+1) OR (4+2+2) OR (4+3+1) OR (3+3+2) \(3+6+3+6+3=\large{21}\) ways to do that. if 1st student gets 1: the others get (8+1+1) OR (7+2+1) OR (5+2+2) OR (5+3+1) OR (4+3+2) OR (4+4+1) OR (3+3+3) \(3+6+3+6+6+3+1=\large{28}\) ways to do that.

OpenStudy (paxpolaris):

Answer: \[1+3+6+10+15+21+28=\Large \color{RED}{84}\]

OpenStudy (paxpolaris):

Another way of doing this: (7+1+1+1) : \(\large4\) ways to distribute like this (6+2+1+1) : \(\large{12}\) ways (5+3+1+1) : \(\large{12}\) ways (5+2+2+1) : \(\large{12}\) ways (4+4+1+1) : \(\large{6}\) ways (4+3+2+1) : \(\large{24}\) ways (4+2+2+2) : \(\large{4}\) ways (3+3+2+2) : \(\large{6}\) ways (3+3+3+1) : \(\large{4}\) ways Total : \(\Large\overline{\color{red}{84}}\) ways